OpenStudy (anonymous):

Given the parabola f(x) = -(x + 2)2 + 4, find the x-intercept(s).

jimthompson5910 (jim_thompson5910):

To find the x-intercepts, plug in f(x) = 0. In other words you replace f(x) with 0. Then you solve for x f(x) = -(x + 2)^2 + 4 0 = -(x + 2)^2 + 4 Now solve for x. Let me know if you still need help

OpenStudy (anonymous):

wait is it x^2+4x+4

jimthompson5910 (jim_thompson5910):

(x + 2)^2 expands out to x^2 + 4x + 4

jimthompson5910 (jim_thompson5910):

So -(x + 2)^2 expands out to -(x^2 + 4x + 4) which distributes to -x^2 - 4x - 4

jimthompson5910 (jim_thompson5910):

This means -(x + 2)^2 + 4 simplifies to -x^2 - 4x

OpenStudy (anonymous):

how do i get my x-ints from that

jimthompson5910 (jim_thompson5910):

We want to solve 0 = -(x + 2)^2 + 4 for x, which is the same as solving 0 = -x^2 - 4x

jimthompson5910 (jim_thompson5910):

Which flips to -x^2 - 4x = 0

jimthompson5910 (jim_thompson5910):

Do you see where to go now?

OpenStudy (anonymous):

sort of

jimthompson5910 (jim_thompson5910):

What do you get when you solve -x^2 - 4x = 0 ?

OpenStudy (anonymous):

thats the thing im not quite sure how to solve it would i get x^2=4x or what

jimthompson5910 (jim_thompson5910):

Factor as much as you can

OpenStudy (anonymous):

or would you take out and x

jimthompson5910 (jim_thompson5910):

so factor out -x to get -x(x + 4) = 0

OpenStudy (anonymous):

right

jimthompson5910 (jim_thompson5910):

then you can break up -x(x + 4) = 0 to get -x = 0 or x + 4 = 0 So you now just need to solve each equation for x

OpenStudy (anonymous):

so you get x=-4 and x=0

jimthompson5910 (jim_thompson5910):

good you got it

jimthompson5910 (jim_thompson5910):

so the x-intercepts as points are (0,0) and (-4,0)

OpenStudy (anonymous):

alright im going to post some more could youi help??

jimthompson5910 (jim_thompson5910):

sure